# Escaping Plato’s Cave: JAM for Aligning Independently Trained Vision and Language Models

### v1

@@ 4 Experiments
We adopt the same data set-up as in the statistical tests, splitting each dataset into 70-15-15 train/validation/test splits. For each task, we train our Joint Autoencoder Modulator (JAM) with Spread loss for 100 epochs with a batch size of 32, using data seeds 5, 42, and 55. Both autoencoders are optimized jointly using AdamW [35] with gradient clipping (1.0) and a cosine annealing scheduler. The reconstruction loss is weighted by a linearly decaying factor $\lambda(t)$, decreasing from 1.0 to 0.1 over training epochs to gradually emphasize the alignment objective. Every five epochs, we compute image-to-text Recall@1 on the validation set, applying early stopping if no improvement is observed for five consecutive validations. We evaluate on two retrieval settings: (1) binary choice between the positive and its hard negative (standard in fine-grained evaluation [19, 28]), and (2) a 5-way choice including three additional distractors. All reported results are averaged across the three seed runs.
To contextualize our approach—designed to escape the “Platonic cave” of unimodal models through post-hoc alignment—we also compare against CLIP [13], a model natively trained for multimodal representation learning. For fairness, we fine-tune CLIP on the same train/val/test splits.
@@ 4.1 Does JAM enable escaping Plato’s Cave?
Our experiment results (Table 2) show that the proposed Spread loss consistently outperforms both the Con and NegCon baseline across retrieval and alignment tasks. This improvement can be attributed to the explicit supervision embedded in Spread, which leverages the notion of “similar context”—a structured grouping based on the similarity of data points. While Con and NegCon optimize for alignment using only pairwise correspondence (and in NegCon’s case, further incorporating explicit hard-negatives), they lack a mechanism for more fine-grained context understanding. Spread, in contrast, directly encourages modality-specific embeddings to reflect not only pairwise similarity, but also semantic proximity across local clusters, resulting in better fine-grained alignment.
Despite its simplicity, our Spread-trained JAM achieves competitive performance not only compared to other contrastive loss variants, but also relative to strong baselines including pretrained and fine-tuned CLIP models. While CLIP benefits from large-scale multimodal training using massive paired datasets and purpose-built architectures, our approach takes an economical path. We do not rely on large-scale image–text corpora or models explicitly designed for multimodal integration. Instead, we build on pretrained unimodal foundations, aligning them post hoc via a lightweight, task-specific joint autoencoding strategy; much cheaper than training both from scratch. This specialist alignment approach sidesteps the need for end-to-end multimodal retraining and architectural entanglement, offering a more economical and modular alternative to generalist training. Our results suggest that meaningful cross-modal structure can emerge through structured alignment—without requiring scale, paired data, or multimodal pretraining.
Refer to caption Refer to caption
@@
To investigate this, we move beyond a fixed-$\alpha$ formulation and adopt curriculum learning [36] as a probing framework (Fig. 5, 5). Curriculum learning [36] is a training paradigm inspired by human learning, where models are first exposed to easier concepts or objectives and gradually introduced to harder or more nuanced distinctions. The motivation is to guide optimization along a smoother path: by focusing early training on easier, more generalizable patterns, the model establishes a robust representational foundation before facing fine-grained or ambiguous cases. We use curriculum learning as a diagnostic framework to probe the nature of supervision required for aligning multimodal embeddings. Specifically, we vary the balancing parameter $\alpha$ in $\mathcal{L}_{spread}$ to control the emphasis between broad semantic alignment (ConCon) and fine-grained contrast (ContextNCE).
@@
By applying this curriculum across embeddings extracted from early, mid, and late layers of both vision and language models, we observe that earlier layers consistently require higher $\alpha$ to achieve optimal alignment — i.e., the epoch where the model achieves the best retrieval accuracy (Fig. 5, 5). This trend aligns with the findings from concept interpretability studies (e.g., TCAV [37], Network Dissection [38]), which show that lower layers tend to represent primitive or entangled features, requiring stronger supervision to isolate meaningful concepts. In our case, increasing $\alpha$ acts analogously — providing finer discriminative signals that help disentangle and align early-layer representations across modalities. In contrast, later-layer embeddings — which are already more abstract and task-aligned — can be aligned effectively with less need for explicit intra-context separation. Notably, we also observe that retrieval performance improves progressively from early to late layers across all task types and model pairings. This suggests that deeper layers not only require less contrastive supervision but also encode more directly alignable representations. Thus, our use of curriculum serves as a controlled lens into the nature of supervision required at different levels of abstraction, offering insight into how independently trained vision and language models organize information across depth.
@@ 4.3 Does JAM performance scale with backbone models’ scale?
To investigate whether increasing the scale of pretrained language backbones improves alignment in our JAM framework, we conduct experiments using three models of varying size: For language models, Gemma2 (9B), LLaMA 3.2 (3B), and OLMo2 (13B), and for vision model, DINOv2 ViT-L. The objective is to understand how representational capacity – where models are trained with larger parameters and embedding dimensions– affects fine-grained multimodal alignment when trained using different contrastive optimization objectives. Gleaned from the results in Table 3, we observe that increasing model scale does not consistently lead to improved performance, regardless of the contrastive method used. The performance of Spread loss remains robust and consistently strong, but the gains from moving to larger-scale language models are minimal or negligible.
One potential reason for this plateau in benefit lies in the data regime. The training set for these spatial reasoning tasks is relatively small, typically in the range of 200 to 1000 image–text pairs. Large pretrained models are known to require substantial fine-tuning data to fully leverage their representational capacity. In our case, the limited data may not provide enough gradient signal for the larger models to differentiate themselves, effectively neutralizing their potential advantage. Another explanation relates to the nature of the task itself. These spatial reasoning benchmarks are fine-grained and often involve small semantic distinctions (e.g., between spatial prepositions or swapped object attributes). Such tasks demand precise local alignment rather than broad generalist knowledge. As a result, the inductive strengths of larger models trained on general web-scale corpora may not translate to meaningful improvements. This reinforces the utility of our task-specific Spread loss, which is explicitly designed to handle such nuanced contrasts.
@@ B.2.3 CCA (Canonical Correlation Analysis)
[…]
@@ B.3.3 Swap task
Gemma2 & DINOv2 0.95 / 0.95 0.317 0.44 0.228 0.47 / 0.47 0.037 0.124 0.014 0.94 / 0.94 0.33 0.433 0.22
Llama3.2 & DINOv2 0.94 / 0.94 0.463 0.473 0.311 0.47 / 0.47 0.049 0.129 0.012 0.94 / 0.94 0.453 0.43 0.287
OLMo2 & DINOv2 0.95 / 0.94 0.466 0.462 0.312 0.48 / 0.48 0.05 0.13 0.012 0.94 / 0.94 0.455 0.511 0.276
@@ C.1 Our Formulation of Spread Loss for Multimodal Representation Learning
$\mathcal{L}_{ConCon}(\Phi_{V},v_{i},B)=-\frac{1}{|C_{L}(i)|}\sum_{c_{l}\in C_{L}(i)}\log\left(\frac{\sigma(v_{i},c_{l})}{\sigma(v_{i},c_{l})+\sum_{\widetilde{c_{l}}\in\widetilde{C_{L}}(i)}\sigma(v_{i},\widetilde{c_{l}})}\right),$ (18)
@@ Appendix D Model Analysis & Training Configuration
D.1 Pretrained State-of-the Art Unimodal Backbones
@@
Table 11: Configuration of pretrained unimodal models used in experiments.
@@ D.2 JAM (Joint Autoencoder Modulator) Framework
Appendix E Compute Systems/Resources

### v5

@@ 4 Experiments
We adopt the same data set-up as in the statistical tests using Sugarcrepe [20], splitting each dataset into 70-15-15 train/validation/test splits. To further test on our method’s performance in different dataset, we incorporated Winoground [34], a benchmark designed to test visuo-linguistic compositionality.
To test on our framework’s versatility across different pretrained backbones, we extracted text embeddings and image embeddings from wider set of models: for language models, Gemma2 (2B,9B) [21], Llama3.2 (1B,3B) [14], OLMo2 (7B,13B) [22], and for vision models, DINOv2 [23], MAE [35], and Swav [36] for self-supervised learning (SSL) objective and ResNet50 [24], Swin [37], and ViT [38] for supervised (Sup) objective. (In Table 2, we show results for DINOv2 and RestNet50 for vision backbones in Sugarcrepe setting. For full results (all possible pretrained backbones configurations and Winoground task), refer to Appendix Table 14, 15.)
For each task, we train our Joint Autoencoder Modulator (JAM) with Spread loss for 100 epochs with a batch size of 32, using data seeds 5, 42, and 55. Both autoencoders are optimized jointly using AdamW [39] with gradient clipping (1.0) and a cosine annealing scheduler. The reconstruction loss is weighted by a linearly decaying factor $\lambda(t)$, decreasing from 1.0 to 0.1 over training epochs to gradually emphasize the alignment objective. Every five epochs, we compute image-to-text Recall@1 on the validation set, applying early stopping if no improvement is observed for five consecutive validations. We evaluate on two retrieval settings: (1) binary choice between the positive and its hard negative (standard in fine-grained evaluation [20, 30]), and (2) 5-way choice including three additional distractors. For Winoground [34], we used their proposed text-score metrics for evaluation, defined as the percentage of instances where, for a fixed caption, the model assigns a higher similarity to the correct image than to the incorrect one. All reported results are averaged across three seed runs.
To contextualize our approach—designed to escape the “Platonic cave” of unimodal models through post-hoc alignment—we compare against CLIP [15], a model natively trained for multimodal representation learning. For fairness, we also fine-tuned CLIP on the same train/val/test splits using the fine-tuning method proposed in [30]. To further motivate our JAM architecture with spread loss, we experimented on simple projection-based methods (linear/non-linear) with our Spread loss as well as JAM with spread loss without reconstruction component (spread w/o reconst.) as ablation experiment (Table 3).
@@ 4.1 Does JAM enable escaping Plato’s Cave?
Language Backbone Vision Backbone Alignment Method Replace Task Add Task Swap Task
Gemma2 (2B) DINOv2 (ViT-B; 86M) Con 65.14 58.34 57.36
NegCon 86.32 96.85 74.5
Spread 88.01 95.74 80.2
Gemma2 (2B) ResNet50 Con 74.23 63.02 63.75
NegCon 87.28 94.80 72.27
Spread 87.60 94.81 82.17
Llama3.2 (1B) DINOv2 (ViT-B) Con 66.97 60.99 66.59
NegCon 83.43 94.25 68.26
Spread 87.21 97.83 79.05
Llama3.2 (1B) ResNet50 Con 67.36 67.94 71.11
NegCon 85.74 94.10 71.69
Spread 87.17 96.85 82.17
OLMo2 (7B) DINOv2 (ViT-B) Con 68.48 65.17 66.88
NegCon 85.30 96.67 71.69
Spread 88.32 97.41 77.62
OLMo2 (7B) ResNet50 Con 68.01 64.35 73.38
NegCon 89.69 90.12 71.38
Spread 86.90 98.44 80.46
Pretrained CLIP (ViT-B-32/OpenAI) [40] 81.05 77.58 64.69
Pretrained CLIP (ViT-B-32/LAION-400m) [40, 41] 80.90 79.64 67.30
Pretrained CLIP (ViT-B-32/SigLIP) [42] 85.01 86.56 70.97
Finetuned CLIP (ViT-B/32) 74.62 92.69 67.21
Table 2: Image-to-Text Retrieval Results of Joint Autoencoder Modulator (JAM) for Vision-Language Compositionality. We report Recall@1 scores (binary) across three compositional tasks in Sugarcrepe: Replace, Add, and Swap, using models trained with different alignment methods in JAM framework. JAM with Spread loss consistently outperforms contrastive baselines across all tasks and backbones. Moreover, it matches or surpasses several strong pretrained and finetuned CLIP variants, highlighting the effectiveness of structured alignment over independently pretrained representations for fine-grained vision–language reasoning.
Post-hoc Alignment Method FLOPs(G) Replace Task Add Task Swap Task
Linear Proj. with Spread 0.06 85.95 91.94 63.33
NonLinear Proj. with Spread 0.12 86.99 92.22 68.57
JAM with Spread w/o reconst. 3.70 83.11 92.53 74.50
JAM with Spread 3.70 88.01 95.74 80.2
Finetuned CLIP (ViT-B/32) 10.48 74.62 92.69 67.21
Table 3: Image-to-Text Retrieval Results in Vision-Language Compositionality with Different Post-hoc Alignment architectures (Linear Projection, Nonlinear Projection, Joint Autoencoder Modulator (JAM)) with Spread Loss and ablation result for removing the reconstruction component in Spread Loss. Same evaluation scheme as shown in Table 2. We show results using Gemma2 (2B) and DINOv2 (ViT-B) as language and vision backbone. We also provide FLOPs(G) across alignment methods (including finetuning CLIP) to show that JAM with Spread framework shows the best performance but requires less FLOPs compared to finetuning.
Our experiment results (Table 2 for Sugarcrepe with binary Recall@1 setting, Table 14 in Appendix for Sugarcrepe with all pretrained backbone configurations with binary & 5-way Recall@1 scores, Table 15 in Appendix for Winoground task) show that the proposed Spread loss consistently outperforms both the Con and NegCon baseline and pretrained/finetuned CLIP across tasks. Spread’s superior performance can be attributed to its structured supervision, which encourages embeddings to reflect local clusters and semantic proximity, thereby enabling a more fine-grained alignment. This contrasts with other methods that rely solely on pairwise correspondence.
Our method achieves this competitive performance using a lightweight, post-hoc alignment approach with a joint autoencoder, which avoids the need for massive paired datasets and architectural complexity associated with pretraining large-scale multiomodal models like CLIP. This makes our method more robust in low-resource or specialized data domains, where fine-tuning a model like CLIP can lead to overfitting (i.e., losing CLIP’s broad priors) and a "representation collapse" [43], as we also observe in our experiments where fine-tuning CLIP led to lower performance 2.
We also conducted stronger baseline experiments using linear and nonlinear projections over frozen embeddings to motivate our JAM framework and ablation experiment of reconstruction component in our spread loss formulation. Note that for projection-based baselines, the reconstruction component of our Spread loss is inherently absent. As Table 3 shows, our method (JAM+Spread) consistently outperformed projection baselines: while projections achieved lower FLOPs, they suffered from significantly lower accuracy and unstable learning due to the lack of reconstruction. Furthermore, in ablation experiments, removing the reconstruction term led to rapid overfitting, with validation loss rising sharply and models requiring early stopping to avoid collapse. This highlights the critical role of reconstruction as a regularizer, preserving generalizable structure across modalities beyond simple alignment.
@@
To investigate this, we move beyond a fixed-$\alpha$ formulation and adopt curriculum learning [44] as a probing framework (Fig. 6, 6). Curriculum learning [44] is a training paradigm inspired by human learning, where models are first exposed to easier concepts or objectives and gradually introduced to harder or more nuanced distinctions. The motivation is to guide optimization along a smoother path: by focusing early training on easier, more generalizable patterns, the model establishes a robust representational foundation before facing fine-grained or ambiguous cases. We use curriculum learning as a diagnostic framework to probe the nature of supervision required for aligning multimodal embeddings. Specifically, we vary the balancing parameter $\alpha$ in $\mathcal{L}_{spread}$ to control the emphasis between broad semantic alignment (ConCon) and fine-grained contrast (ContextNCE).
@@
By applying our curriculum across embeddings from early, mid, and late layers of vision and language models, we find that earlier layers consistently require higher $\alpha$ for optimal alignment—the point of peak retrieval accuracy (Fig. 6, 6). This mirrors interpretability studies (e.g., TCAV [45], Network Dissection [46]) showing that lower layers encode primitive or entangled features, demanding stronger supervision to isolate meaningful concepts. Here, larger $\alpha$ provides the finer discriminative signal needed to disentangle and align these early representations. Later layers, already more abstract and task-aligned, align effectively with weaker supervision. Retrieval performance also improves steadily from early to late layers, indicating that deeper layers both require less contrastive pressure and yield more directly alignable representations. Thus, the curriculum offers a controlled lens into the supervision needed across representational depth, revealing how independently trained vision and language models organize information.
@@ 4.3 Does JAM performance scale with backbone models’ scale?
Language Backbone Vision Backbone Alignment Method Replace Task Add Task Swap Task
Gemma2 (9B) DINOv2 (ViT-L; 300M) Con 68.77 59.31 60.92
NegCon 87.52 98.89 84.44
Spread 89.66 98.89 84.44
Llama3.2 (3B) DINOv2 (ViT-L) Con 65.24 62.38 73.12
NegCon 84.54 92.61 75.93
Spread 89.43 95.74 82.44
OLMo2 (13B) DINOv2 (ViT-L) Con 65.86 59.21 62.28
NegCon 89.36 93.7 78.89
Spread 90.53 97.96 84.71
Table 4: Model Scaling Experiment. We report image-to-text Recall@1 scores across three spatial reasoning tasks (Replace, Add, Swap) using language backbones of increasing scale—Gemma2 (9B), LLaMA 3.2 (3B), and OLMo2 (13B)—within our JAM framework. Despite varying parameter counts, we observe no consistent trend of performance improvement with model size. Spread loss performs robustly across all scales, suggesting that for fine-grained, low-data tasks, representational alignment depends more on objective design than on model scale.
We test whether larger pretrained language backbones improve alignment in JAM, using Gemma2 (9B), LLaMA 3.2 (3B), and OLMo2 (13B) with a fixed vision encoder (DINOv2 ViT-L). The aim is to assess how representational capacity affects fine-grained multimodal alignment under different contrastive objectives.
As shown in Table 4, scaling language models does not consistently improve performance. Spread loss remains robust, but gains from larger models are minimal. This plateau likely reflects the small training regime (200–1000 image–text pairs), which limits the gradient signal needed for large models to leverage their capacity. Moreover, the fine-grained nature of spatial reasoning tasks emphasizes precise local alignment (e.g., prepositions, swapped attributes), reducing the advantage of broad generalist knowledge from web-scale pretraining. These results highlight the effectiveness of task-specific Spread loss in capturing nuanced multimodal contrasts.
@@ B.2.3 CCA (Canonical Correlation Analysis)
$\text{s.t.}\quad\mathbf{w}^{\!\top}\Sigma_{xx}\mathbf{w}=1,\;\mathbf{v}^{\!\top}\Sigma_{yy}\mathbf{v}=1.$ (13)
@@ B.3.3 Swap task
Gemma2 & DINOv2 0.95 / 0.95 0.317 0.440 0.228 0.47 / 0.47 0.037 0.124 0.014 0.94 / 0.94 0.330 0.433 0.220
Llama3.2 & DINOv2 0.94 / 0.94 0.463 0.473 0.311 0.47 / 0.47 0.049 0.129 0.012 0.94 / 0.94 0.453 0.430 0.287
OLMo2 & DINOv2 0.95 / 0.94 0.466 0.462 0.312 0.480 / 0.48 0.05 0.130 0.012 0.94 / 0.94 0.455 0.511 0.276
Gemma2 & ResNet50 0.89 / 0.89 0.315 0.430 0.221 0.50 / 0.50 0.029 0.123 0.014 0.89 / 0.89 0.315 0.428 0.220
@@ C.1 Our Formulation of Spread Loss for Multimodal Representation Learning
[…]
@@ Appendix D Model Analysis & Training Configuration
D.1 Pretrained Unimodal Backbones
@@
MAE (SSL) 86M 12 768
Swav (SSL) 23M 50 2048
Swin (Sup) 88M 50 2048
ViT (Sup) 86M 12 768
ResNet50 (Sup) 23M 50 2048
Table 12: Configuration of pretrained unimodal models used in experiments. For MAE, we used ViT as backbone architecture. For Swav and Swin, we used RestNet50 as backbone architecture.
@@ D.2 JAM (Joint Autoencoder Modulator) Framework
Appendix E Further Results
Language Backbone Vision Backbone Alignment Method Replace Task Add Task Swap Task
(Model Size) (Model Size) for JAM Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way)
Gemma2 (2B) MAE (SSL) Con 58.55 50.33 56.37 50.46 63.73 49.29
NegCon 77.62 69.38 89.15 80.16 70.81 58.63
Spread 89.09 71.16 94.35 84.98 79.05 58.63
Gemma2 (2B) Swav (SSL) Con 68.08 62.34 65.25 54.41 62.60 59.21
NegCon 83.31 76.53 95.20 91.63 80.75 67.04
Spread 88.88 76.98 95.83 90.64 81.01 67.70
Gemma2 (2B) Swin (Sup) Con 64.94 58.17 56.87 50.66 62.91 59.52
NegCon 85.07 78.35 96.39 91.66 77.93 70.92
Spread 85.16 79.01 97.87 92.22 80.32 71.46
Gemma2 (2B) ViT (Sup) Con 60.78 54.74 60.79 55.05 59.37 52.28
NegCon 86.06 75.89 96.31 92.50 59.21 49.58
Spread 86.35 75.26 95.37 89.62 82.31 67.72
NegCon 87.52 84.33 98.89 93.14 84.44 69.69
Spread 89.66 83.34 98.89 96.30 84.44 73.38
Llama3.2 (1B) MAE (SSL) Con 67.82 54.48 61.08 55.33 61.19 54.68
NegCon 82.92 74.09 94.16 84.80 76.77 68.26
Spread 86.23 74.96 96.85 92.03 80.47 68.67
Llama3.2 (1B) Swav (SSL) Con 65.94 58.97 68.12 62.93 75.39 68.03
NegCon 92.37 83.91 92.12 87.48 78.62 71.84
Spread 88.63 79.24 98.94 94.75 83.29 69.99
Llama3.2 (1B) Swin (Sup) Con 70.59 63.60 69.70 63.48 75.39 68.03
NegCon 86.01 83.05 93.89 88.39 69.11 63.45
Spread 87.24 76.69 94.72 94.16 77.62 68.84
Llama3.2 (1B) ViT (Sup) Con 66.81 57.49 64.50 60.70 64.87 59.79
NegCon 82.81 77.15 93.16 90.57 74.50 63.18
Spread 87.56 76.13 97.38 92.23 74.23 57.79
Llama3.2 (3B) DINOv2 (ViT-L) Con 65.24 59.14 62.38 57.9 73.12 65.18
NegCon 84.54 82.66 92.61 88.16 75.93 70.27
Spread 89.43 82.48 95.74 91.25 82.44 76.77
OLMo2 (7B) MAE (SSL) Con 66.88 57.37 68.12 56.98 58.36 51.85
NegCon 86.39 74.57 93.05 88.33 71.94 60.62
Spread 83.05 77.84 96.39 90.64 79.32 64.52
OLMo2 (7B) Swav (SSL) Con 66.38 61.12 67.38 62.58 63.75 58.65
NegCon 88.18 83.31 92.12 88.50 82.23 69.98
Spread 93.27 85.63 98.43 93.70 81.01 72.81
OLMo2 (7B) Swin (Sup) Con 65.35 59.45 70.82 63.79 55.67 50.85
NegCon 89.40 86.70 95.28 89.35 73.66 71.96
Spread 91.32 82.11 97.87 94.72 80.16 70.81
OLMo2 (7B) ViT (Sup) Con 70.56 64.79 69.98 58.86 69.15 60.36
NegCon 84.61 80.77 94.16 87.86 81.74 67.30
Spread 89.25 81.01 98.43 95.18 76.49 66.44
OLMo2 (13B) DINOv2 (ViT-L) Con 65.86 68.68 59.21 62.52 62.28 52.12
NegCon 89.36 82.68 93.7 91.66 78.89 67.57
Spread 90.53 83.18 97.96 92.42 84.71 69.69
Table 14: Image-to-Text Retrieval Results of Joint Autoencoder Modulator (JAM) for Vision-Language Compositionality with wider set of pretrained backbones.
Language Backbone Vision Backbone Winoground Text-score
Gemma2 (2B) DINOv2 (ViT-B; 86M) 58.75
ResNet50 58.13
Llama3.2 (1B) DINOv2 (ViT-B) 57.5
ResNet50 61.3
OLMo2 (7B) DINOv2 (ViT-B) 55
ResNet50 57.5
Pretrained CLIP (ViT-B/32) 32.80
Table 15: Winoground Text-score results using JAM with Spread framework. Given the observation from Sugarcrepe in Table 2, 3 that finetuning CLIP in small data regime leading to decrease in performance, we provide pretrained CLIP as a baseline. We also observe JAM with Spread method showing the best performance compared to the baseline.
Appendix F Compute Systems/Resources